### v1

# Squeezing information from radio surveys to probe the primordial Universe

Abstract
A major goal of cosmology is to understand the nature of the field(s) which drove primordial Inflation. Through future observations, the statistics of large-scale structure will allow us to probe primordial non-Gaussianity of the curvature perturbation at the end of Inflation. We show how a new correlation statistic can significantly improve these constraints over conventional methods. Next-generation radio telescope arrays are under construction which will map the density field of neutral hydrogen to high redshifts. These telescopes can operate as an interferometer, able to probe small scales, or as a collection of single dishes, combining signals to map the large scales. We show how to fuse these these operating modes in order to measure the squeezed bispectrum with higher precision and greater economy. This leads to constraints on primordial non-Gaussianity that will improve on measurements by Planck, and out-perform other surveys such as Euclid. We forecast that $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\sim 3$, achieved by using a small subset, $\mathcal{O}(10^{2}-10^{3})$, of the total number of accessible triangles. The proposed method can be used for any signal that peaks in squeezed configurations.
I Introduction
Primordial non-Gaussianity (PNG) is one of the most promising observational avenues to probe the primordial Universe, due to its sensitivity to the physics of these early stages. Measuring PNG could reveal the field content and the amplitude of field interactions [1], as well as verify the quantum nature of the primordial fluctuations [2] that seed the large-scale structure (LSS) we observe today, as predicted by Inflation. This early Universe information can be extracted by characterizing how much the distribution of the primordial perturbations deviates from Gaussianity.
We propose a novel method that utilises the functionalities of a radio telescope array in 21cm intensity mapping surveys, in order to measure the bispectrum of squeezed triangles, in an optimal and efficient way, while maximizing the PNG signal and avoiding regions contaminated by systematics. The squeezed bispectrum, with $k_{1}\approx k_{2}\gg\,k_{3}$, can be used to significantly improve current bounds on local PNG [3, 4, 5, 6]. Constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$ provides a unique observational handle to rule out entire classes of inflationary models and enhance our knowledge of the early Universe [7].
Currently, the tightest constraints on PNG are from cosmic microwave background (CMB) measurements [8]. The results imply that the primordial perturbation field is weakly non-Gaussian and hence most of the PNG information is contained in the primordial bispectrum, i.e. the lowest-order deviation from a Gaussian distribution. Most of the available information on local PNG from the CMB anisotropies bispectrum has been extracted. LSS surveys have the potential to significantly improve the current PNG bounds [9, 10], since a 3D dark matter tracer (e.g. galaxies) has considerably more bispectrum modes than the 2D CMB. However, most of these LSS modes are in the nonlinear regime, which is dominated by the gravitational evolution of cosmic structures. The subdominant PNG signal must be separated from the late-time non-Gaussian gravitational contributions with significant precision. Recent results from optical galaxy surveys [11, 12, 13, 14, 15], where an analytical perturbative model has been used (see e.g. [16]), are far from being competitive with the CMB bounds. The main reason is the complexity of accurate analytical models beyond the perturbative regime which is needed to extract the wealth of PNG information from the nonlinear scales. A promising alternative is to rely on numerical approaches, like simulation-based [17, 18, 19, 20] or field-level inference [21, 22].
In this letter, we employ a perturbative analysis at high redshifts, which facilitates access to a large number of bispectrum modes in the perturbative regime. By also using high-density samples, we can significantly boost the PNG signal. To achieve this, we investigate upcoming high-redshift 21cm intensity mapping surveys [23, 24, 25, 26, 27, 28, 29, 30], focusing on the squeezed bispectrum, $k_{1}\approx k_{2}\gg\,k_{3}$. Due to its distinct triangle preference, which couples different scales, this shape is mostly orthogonal to the late-time gravitational part of the LSS bispectrum and therefore relatively easier to separate from the latter.
II PNG from 21cm intensity mapping
In the post-reionization Universe most of the neutral hydrogen (HI) resides inside galaxies, thus detecting its 21cm emission line provides a tracer of the underlying matter distribution. The HI intensity mapping (IM) technique consists of making brightness temperature maps of the sky at different radio frequencies, by measuring the integrated emission line from many galaxies that reside within each pixel. The combined emission yields a large detectable signal, while probing very large volumes of the Universe from the epoch of re-ionisation up to present times. Such spectroscopic surveys are planned for example with the Square Kilometre Array Observatory (SKAO) [29] and the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX) [30] and will cover a wide sky area and deep redshift range ($0<z\lesssim 3$).
An array of radio telescopes are set up to measure HI intensity in two distinct ways: in single-dish (SD) mode, by auto-correlating the signal of each dish individually or in interferometer mode (IF), where the signal from all dishes is cross-correlated. The first probes the large and linear scales of clustering, while the latter probes the intermediate to the very small and nonlinear scales (large $k$-mode values) with high angular resolution. Both modes of the instrument compliment each other in Fourier space, as shown in the schematic illustration of Fig. 1. Additionally the two functionalities differ in the behaviour of the instrumental noise, which is the dominant noise contribution [31]. Contrary to optical surveys, the shot noise can be negligible, up to small scales.
[…]
Figure 1: Schematic illustration of the ranges of radial and transverse wavenumbers that the two survey modes are sensitive to (adapted from [43]).
The presence of local PNG induces a large-scale scale-dependent behaviour on the power spectrum of a tracer [44, 45, 46, 47, 48, 49]. Hence, a LSS survey should be able to probe the largest scales of clustering in order to provide competitive constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$. Moreover, attaining a deeper comprehension of galaxy formation is essential for establishing constraints on the amplitude of local PNG [50, 51, 52]. For the bispectrum, access to a wide scale range is essential, in order to probe enough squeezed triangles and boost the signal-to-noise ratio. Thus, a single-dish HI IM survey favours a power spectrum analysis, where has been suggested as well in the past as a powerful probe for local PNG [53, 54, 55, 56, 57], outperforming an interferometer survey. The reverse holds for the bispectrum $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, where surveys designed for IF mode outshine those in SD mode [58].
Due to the intrinsic instrumental and observational scale restrictions, HI IM surveys have not yet reached their full potential in constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$, despite their optimal specifications (i.e. wide sky area, high redshifts and sample densities). This is even more evident for the bispectrum of an IF survey, where a more futuristic setup, like PUMA [59] or SKA2 LOW [60] is required to significantly improve over current bounds [61, 58] .
III The 21cm bimodal bispectrum
In this Letter, we propose a novel method to push the capabilities of HI IM to tightly constrain the amplitude of local PNG, while excluding the contaminated $k$-modes from the analysis. The key observation lies in the fact that the two functionalities of the radio telescope probe complimentary scales (see Fig. 1). Combining the data from these two can provide access to a very wide scale range with high resolution: from the large and linear scales, as probed by SD mode, up to the very small and nonlinear ones, measured by IF mode. This is optimal for a bispectrum analysis and especially for an observable that has a strong signal on squeezed triangles, like local PNG.
Squeezed configurations correlate different scales, i.e. one side of the triangle is much smaller than the other two ($k_{1}\sim k_{2}\gg k_{3}$). Thus, cross correlating a density field $\delta^{\rm SD}(\bm{k})$, as probed by a single-dish mode survey, with two density fields $\delta^{\rm IF}(\bm{k})$ probed by a survey in interferometer mode, we can form an HI IM bispectrum that has access to a very large number of squeezed triangles. We define the bimodal bispectrum, synergising two HI IM surveys in two different operating modes, for an overlapping redshift range and sky area, as
$\langle\delta^{\rm IF}(\bm{k}_{1})\,\delta^{\rm IF}(\bm{k}_{2})\,\delta^{\rm SD}(\bm{k}_{3})\rangle=\\ \,(2\,\pi)^{3}\,\delta_{\rm D}(\bm{k}_{123})\,B^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\;,$ (1)
where $\bm{k}_{123}\equiv\bm{k}_{1}+\bm{k}_{2}+\bm{k}_{3}$ and the Dirac $\delta_{\rm D}$ ensures momentum conservation.
The fields $\delta^{\rm SD,IF}$ are fluctuations in the 21cm brightness temperature and their measurements in Fourier space depend on survey window functions $W^{\rm SD,IF}(\bm{k})$. Crucially, SD observations observe the actual sky, so that the fluctuation is a convolution in Fourier space, ${\delta^{\rm SD}(\bm{k})=W^{\rm SD}(\bm{k})\star\delta(\bm{k})}$, whereas IF mode measures a Fourier transform of the sky, leading to a simple product, ${\delta^{\rm IF}(\bm{k})=W^{\rm IF}(\bm{k})\,\delta(\bm{k})}$. Here and below we omit the $z$-dependence for brevity.
[…]
The estimator for measuring the synergy bispectrum between the SD and IF modes of the neutral hydrogen instrument is
$\hat{B}^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\frac{1}{V_{s}\,V_{123}}\sum_{\bm{q}_{1}\in k_{1}}\sum_{\bm{q}_{2}\in k_{2}}\sum_{\bm{q}_{3}\in k_{3}}\delta_{\rm K}(\bm{q}_{123})$
×$\times\,\delta^{\rm IF}(\bm{q}_{1})\,\delta^{\rm IF}(\bm{q}_{2})\,\delta^{\rm SD}(\bm{q}_{3})\;,$ (2)
where $V_{s}$ is the volume of the survey within each redshift bin. The Kronecker delta ensures the formation of ‘fundamental triangles’ with sides $\bm{q}_{i}$, satisfying $\bm{q}_{123}=0$, that fall in the ‘triangular bin’ of width $\Delta k$ defined by the bin centres ($k_{1}$, $k_{2}$, $k_{3}$). The volume in Fourier space of the fundamental triangle bins is, in the thin shell limit $V_{123}=8\,\pi^{2}\,k_{1}\,k_{2}\,k_{3}\,\Delta k^{3}$ [67], where the bin size $\Delta k$ is taken to be the fundamental frequency of the survey, $\Delta k=k_{\rm f}=2\pi/V_{\rm s}^{1/3}$. For the extreme cases of flattened ($k_{1}=k_{2}+k_{3}$) and open triangles ($k_{1}\neq k_{2}+k_{3}$) this expression breaks. Here we consider the analytic corrections for these configurations presented in [68].
The covariance of the bimodal bispectrum estimator is given by
${\sf C}_{ij}\equiv\big{\langle}\delta\hat{B}_{i}\,\delta\hat{B}_{j}\big{\rangle}={\sf C}_{ij}^{\rm G}+{\sf C}_{ij}^{\rm NG}\,,$ (3)
where $\delta\hat{B}_{i}=\hat{B}_{i}-\langle\hat{B}_{i}\rangle$, $i$ and $j$ indices denote the different triangle configuration, i.e. $B_{i}\equiv B(k_{1}^{i},k_{2}^{i},k_{3}^{i})$, and G, NG are the Gaussian and non-Gaussian contributions. In the thin shell limit ($k\gg\Delta k$), the Gaussian part can be written as
${\sf C}_{ij}^{\rm G}=\,\frac{(2\pi)^{6}}{V_{s}\,V_{123}}\,s_{123}\,\delta_{ij}\,P_{\rm HI}^{\rm IF}(\bm{k}_{1})\,P_{\rm HI}^{\rm IF}(\bm{k}_{2})\,P_{\rm HI}^{\rm SD}(\bm{k}_{3})\,,$ (4)
where $s_{123}=6,2,1$ for equilateral, isosceles and scalene triangles respectively. The HI IM power spectrum is the tree-level model, including the instrumental noise and shot noise [58] (see also the Appendix).
The non-Gaussian part of the covariance is a complicated quantity to calculate analytically. However, in the case of squeezed configurations, like those considered in this work, it can be accurately approximated by [69, 68, 70]
${\sf C}_{ij}^{\rm NG}$ $=\frac{2(2\,\pi)^{3}}{V_{s}\,V_{123}^{i}\,V_{123}^{j}}\Big{[}\delta_{k_{1}^{i}k_{1}^{\,j}}\,U(k_{1}^{i},k_{1}^{\,j})$
×$\times B_{\rm HI}(\bm{k}_{1}^{\,j},\bm{k}_{2}^{i},\bm{k}_{3}^{i})\,B_{\rm HI}(\bm{k}_{1}^{i},\bm{k}_{2}^{\,j},\bm{k}_{3}^{\,j})+8\,{\rm perm}\Big{]},$ (5)
where $U(k_{1}^{i},k_{1}^{j})=16\pi^{3}\,k_{2}^{i}\,k_{3}^{i}\,k_{2}^{\,j}\,k_{3}^{\,j}\Delta k^{5}$. For the extreme triangle configurations, we consider the analytic corrections presented in Ref. [68]. For the HI IM bispectrum, $B_{\rm HI}$, the tree-level model is considered [58] (see also the Appendix). A well motivated assumption for the instrumental noise is that it is Gaussian [43]; in this case the non-Gaussian part of the covariance will not contain instrumental noise contributions.
IV Results
In this section, we present forecasts for the HI IM bimodal bispectrum on constraining the amplitude of local PNG. The Fisher information matrix formalism is used, which provides the minimum error on a parameter achieved by a survey setup, while requiring only an analytic model as an input. In a given redshift bin the Fisher matrix is
$F_{\alpha\beta}=\int{\rm d}^{3}\bm{k}_{i}\,\frac{\partial B^{\rm SD\times IF}(\bm{k}_{i})}{\partial\theta_{\alpha}}\,{\sf C}_{ij}^{-1}\,\frac{\partial B^{\rm SD\times IF}(\bm{k}_{j})}{\partial\theta_{\beta}}\,.$ (6)
[…]
Here we only consider a squeezed subset of these configurations, i.e. those that satisfy $k_{1}\approx k_{2}\geq 3\,k_{\rm 3}$, since they contain most of the signal for the local PNG. Note, that the bimodal bispectrum can benefit all triangle shapes (see the Appendix). The minimum value is $k_{\rm min}=k_{\rm f}$, which is the largest scale probed in the redshift bin by the survey, and the maximum value $k_{\rm max}$ corresponds to the smallest scale where the theoretical model is reliable. We use a conservative choice $k_{\rm max}(z)=0.75\,k_{\rm NL}(z)$ and an optimistic one $k_{\rm max}(z)=k_{\rm NL}(z)$, where $k_{\text{NL}}$ is given by the inverse square root of the one-dimensional velocity dispersion [61]. The analysis is thus confined within the perturbative regime, where the tree-level model offers good agreement with the numerical results [72, 73, 74, 75, 76]. The tree-level description is used to model the expected value of the bimodal estimator [Section III], including the observational windows for the two operating modes of the instrument. Any $\bm{k}$-mode outside of the window will result in the exclusion of the triangle from the Fisher matrix calculation.
Figure 2: Marginalized forecasts of $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})$ from the bimodal bispectrum (SKA $\times$ HIRAX) with and without foreground cut, for pessimistic and optimistic $k_{\rm max}$. Forecasts from each survey and from the summed signal without cross correlation (SKA + HIRAX) are also shown. Purple line is the Planck 2018 constraint [77].
The forecast error on $\theta_{\alpha}$ is then given by $\sigma(\theta_{\alpha})=(\tilde{F}_{\alpha\alpha}^{-1})^{1/2}$, where $\tilde{F}_{\alpha\beta}=\Sigma_{i}F_{\alpha\beta}(z_{i})$ encompasses the total information from all redshift bins considered. All the parameters that control the model are considered free and are entries of the Fisher matrix. These include the $\Lambda$CDM cosmology parameters and the HI IM bias parameters [9, 61, 58, 78] (see the Appendix, and [79] for further details on the Fisher formalism).
All work on constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$ is affected by a degeneracy between $f_{\mathrm{NL}}^{\mathrm{loc}}$ and $b_{\phi}$ (see e.g. [80, 51]), where $b_{\phi}$ is the local PNG bias parameter [48, 81]. This compromises the ability of the LSS power spectrum and bispectrum to tightly constrain $f_{\mathrm{NL}}^{\mathrm{loc}}$. PNG bias depends on the halo formation history [82], i.e., on properties beyond total mass. If it is accurately modeled and constrained, e.g. with hydrodynamical simulations, this degeneracy is broken and the power of the LSS bispectrum to constrain $f_{\mathrm{NL}}^{\mathrm{loc}}$ is completely restored [52]. Our focus here is on the relative improvement from a new bimodal synergy, compared to the sum of the standard bispectrum information from 2 surveys – therefore, our results are not affected by the model for $b_{\phi}$, and we use the simplest model that breaks the degeneracy [44, 46, 45].
After marginalization over the free parameters of the model [58, 78] (see also the Appendix), we retrieve the final forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$ from the bimodal bispectrum. The two surveys that we consider are Band 1 of SKA-MID [29], which is a radio telescope in SD mode and has a redshift range of $z=0.35-3.05$, and the final version of HIRAX [30] in IF mode with $z=0.75-2.55$. In the analysis, we consider only the signal from the overlapping range ($z=0.75-2.55$) and sky coverage ($15,000$ $\rm deg^{2}$). For details of the surveys considered and the specifics of the instrumental noise, see [78]. Note that the same survey can be used, operating in SD and IF mode, to produce the bimodal synergy. However, telescope design is usually optimized for one of the modes and is not effective in the other.
The forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$ from the synergy between SKA and HIRAX are presented in Fig. 2, for the pessimistic and optimistic $k_{\rm max}$ cases. The results when removing the foreground cut, i.e. when assuming perfect foreground cleaning and signal reconstruction, are shown in order to illustrate the limiting capabilities of the surveys. We also present the forecasts from each individual survey, as well as their summed signal (SKA+HIRAX) without cross-correlation, for the same redshift range and squeezed configurations as in the bimodal case. This allows for a fair comparison and an assessment of the gain from the proposed method. The results indicate that the bimodal bispectrum provides a significant improvement on $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, more than a factor of $3$, compared to those from each individual survey, as well as their summed signal (a factor of $\sim 3$).
The constraints are competitive with current CMB bounds, exhibiting a factor of $\sim 2$ improvement. The bimodal bispectrum achieves this by only using a small subset, $\mathcal{O}(10^{2}-10^{3})$ triangles, of the total number of formed triangles. This means that the forecasts could improve further once all the triangles are considered, even achieving $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\lesssim 1$, as in the case of dedicated surveys (e.g. SPHEREx [83] or PUMA [84]). In addition, by including the single-survey information from the non-overlapping redshifts and sky area, the constraints would improve even further and utilise to the fullest the available signal in the SKA-HIRAX synergy.
V Conclusions
The MeerKLASS pilot survey [85] on the MeerKAT precursor telescope of SKA-MID, has provided HI IM maps both in SD and IF mode. This has allowed the detection of the SD power spectrum in cross correlation with galaxies [86] and the auto-power in IF mode [87]. The measurements in SD mode will see a major advance when the SKA is operating, while the early IF mode measurements will be vastly improved when HIRAX takes data.
We have shown how to combine these observations in a novel way to give significantly stronger constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$ than was previously thought possible. In fact, this method will improve on the Planck results by around a factor of 2, using only a few thousand triangles per redshift bin. The bimodal bispectrum improves constraints relative to the sum of the individual surveys by a factor of $\sim 4$, which is equivalent to a much larger SD or IF array. It is in fact the only way to get any strong constraints on local PNG from the HI intensity bispectrum.
We highlight the point that a substantial part of the constraining power comes from the IF mode signal. Therefore, the results will improve even further if the IF instrumental noise can be reduced. By contrast, the method is not so sensitive to the SD mode noise (see Fig. 4). Instead, the SD signal is sensitive to the loss of very large scales $k_{\parallel}<k_{\parallel{\rm min}}$ (see Figs. 3 and 5), so that further improvement requires advances in foreground cleaning and long-wavelength $k_{\parallel}$-mode reconstruction.
Long-wavelength modes provide a set of very squeezed triangles that carry a significant amount of signal on local PNG. On the other hand, increasing the number of squeezed triangles by changing $k_{\rm max}$ (i.e. the pessimistic and optimistic cases) provides a much less pronounced gain in the $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints ($\sim 20\%$). Notably, the constraints with an optimistic $k_{\rm max}$ are very close ($\sim 15\%$ difference) to those provided by the pessimistic case, in which no radial foregrounds are considered (see Fig. 2). This indicates that the local PNG signal lost to foregrounds can be partially recovered, by venturing further into the intermediate regime. Smaller scales have a significant amount of information on PNG, which requires alternative methods to the analytical modelling, e.g. simulated-based inference [17, 18, 19, 20]. We leave this for future work.
The bimodal bispectrum cross-correlates SD large scales with smaller-scale IF modes. We do not expect foregrounds or instrumental systematics to show correlations between long- and short-wavelength $k$-modes. This is another advantage of the bimodal synergy. The advantages of the bimodal bispectrum are not limited to local PNG studies. Any observable that has signal peaking on the squeezed triangles can benefit from the proposed synergy. Moreover, including all shapes of triangles into the bimodal analysis (see Fig. 6), still maintains a significant gain in the constraining power.
In the case of bispectrum probes of optical surveys, the data are observed in real space, thus a Fourier transformation takes place from the observed three-point correlation function, including the window functions for each side of the triangle. This introduces complex convolutions between the windows and significantly hinders a bispectrum analysis. In the case of the bimodal synergy bispectrum, the IF mode measures a Fourier transformation of the sky, while the SD mode the actual sky, including the observational windows. This means that the window convolutions will be much more under control, offering an additional benefit to the proposed approach.
Acknowledgements DK, RM and SS acknowledge support for this work from the University of Missouri South African Education Program. RM is supported by the South African Radio Observatory and National Research Foundation (grant no. 75415). SS acknowledges support for this work from NSF-2219212. SS is supported in part by World Premier International Research Center Initiative, MEXT, Japan. SC acknowledges support from the Italian Ministry of University and Research, PRIN 2022 ‘EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe Research’, from the Italian Ministry of Foreign Affairs and International Cooperation (grant no. ZA23GR03), and from the European Union – Next Generation EU. JF acknowledges support of Fundação para a Ciência e a Tecnologia through the Investigador FCT Contract No. 2020.02633.CEECIND/CP1631/CT0002, the FCT project PTDC/FIS-AST/0054/2021, and the research grants UIDB/04434/2020 and UIDP/04434/2020.